The ball is tied to a string to whirl it having a radius of 50 cm at a velocity of 1.2 m/s. What will be the acceleration of the ball?

1
Expert's answer
2021-11-01T12:53:37-0400

We can find the acceleration of the ball as follows:


"a_c=\dfrac{v^2}{r}=\dfrac{(1.2\ \dfrac{m}{s})^2}{0.5\ m}=2.88\ \dfrac{m}{s^2}."